OpenStudy (anonymous):

Estimate the square root to the nearest integer. Find two integers that make the equation y^2 = 25 true.

OpenStudy (anonymous):

A. 5, -5 B. 5, 0.5 C. \[-\sqrt{5}, \sqrt{5}\] D. 5, 25

OpenStudy (kinggeorge):

For this one, just try some values. Does \(5^2=25\)? What about \((-5)^2\)?

Parth (parthkohli):

See, there are always two square roots for any number. One is its real square root and the second one is its additive inverse. :D

OpenStudy (anonymous):

so the additive inverse is not a real root? Is it a false root?

Parth (parthkohli):

@FoolForMath Real not as in real number, I mean as its positive square root.

OpenStudy (kinggeorge):

@ParthKohli If the roots are complex, the additive inverse is not necessarily the second root.

OpenStudy (anonymous):

i think it was D right

OpenStudy (kinggeorge):

What is \(25^2\)?

OpenStudy (anonymous):

625

OpenStudy (kinggeorge):

So would 25 satisfy the equation?

Parth (parthkohli):

@KingGeorge I did mention that. See it above.

Parth (parthkohli):

Hint, one is a positive and one is a negative.

OpenStudy (anonymous):

Okay, so you mean that -5 is not real?

OpenStudy (anonymous):

so @ParthKohli it was A

Parth (parthkohli):

No, no, no. I mean that if it has a real square root, then the other square root would be its additive inverse.

Parth (parthkohli):

It is @ZhangYan

OpenStudy (anonymous):

@ZhangYan: Yes it is A, I think you will find this interesting \[ \large \sqrt[n]{a^n} = a \text{ if $n$ is odd } \]\[ \large \sqrt[n]{a^n} = |a| \text{ if $n$ is even } \]

OpenStudy (anonymous):

oh that make sence thanks ffm and parthkohli

OpenStudy (anonymous):

Glad to help :)